OpenStudy (anonymous):

how would you find c in terms of k for this equation:

OpenStudy (anonymous):

\[2^c=\sqrt{\frac{ 4 }{ 2^k }}\]

OpenStudy (anonymous):

id probably say take log of both sides :)

OpenStudy (anonymous):

I haven't done log yet :/

OpenStudy (anonymous):

well that stinks :(

OpenStudy (anonymous):

ok, so make the base of 2 sides equal...then equate the powers :)

OpenStudy (aroyni):

yes take log base 2 \[c=\sqrt(\frac{4}{2^k})\]

OpenStudy (aroyni):

sorry

OpenStudy (aroyni):

i missed the log base two

OpenStudy (anonymous):

I'm not sure how to write \[\sqrt{\frac{ 4 }{ 2^k }}\] as 2 to the power of something...

OpenStudy (anonymous):

ok lets work on that Zara

OpenStudy (aroyni):

\[c=\log_2(\sqrt{\frac{4}{2^k}})\]

OpenStudy (aroyni):

sorry for interupting

OpenStudy (anonymous):

\[\frac{4}{2^k}=\frac{2^2}{2^k}=2^{2-k}\]right?

OpenStudy (anonymous):

aroyni - i don't know what log is, sorry

OpenStudy (anonymous):

ohhh yes i get it now, thank you!

OpenStudy (anonymous):

yw :)

OpenStudy (aroyni):

just don't forget the square root adds a 1/2 so c=1-k/2

OpenStudy (anonymous):

yep, thank you :)